OpenStudy (anonymous):

Eugene wants to buy jeans at a store that is giving $10 off everything. The tag on the jeans is marked 50% off. The original price is $49.98. a. Find the total cost if the 50% discount is applied before the $10 discount. b. Find the total cost if the $10 discount is applied before the 50% discount.

OpenStudy (anonymous):

A) We can create a formula to represent the situation, where "p" equals price: .5(49.98) - 10 = p When we evaluate, we get: (24.99) - 10 = p 14.99 = p B) We can create a new formula, where "p" still equals price: .5(49.98 - 10) = p Evaluate once again: .5(39.98) = p 19.99 = p Hope I helped! Let me know if you need any more help or have any other questions :)

OpenStudy (anonymous):

Thank you soooo much, this helped me a lot!

OpenStudy (anonymous):

Anytime, glad I could help! :D

OpenStudy (anonymous):

what is the total cost if the 50% discount is applied before the $10 discount?

563blackghost (563blackghost):

We would first divide .5 which is the decimal form of 50% with the original price... \[49.98 \times .5 = discount|price\] then you would get your answer in which you then subtract 10 from and you will get your answer ^^

OpenStudy (anonymous):

Thank you this is a great answer! Thank you soooo much!

563blackghost (563blackghost):

ooo meant multiply not divide

OpenStudy (anonymous):

oh, ok thank you anyways!!!!!

563blackghost (563blackghost):

Your welcome ^^ and.... \(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (anonymous):

LOL THANK YOU!

563blackghost (563blackghost):

^^